OpenStudy (anonymous):

Can someone just help me solve this please?? i need an extraneous solution and it solved. i chose this but im not sure its correct. 2√x+5+7=4

OpenStudy (stefrheart):

Is that the original equation?

OpenStudy (anonymous):

no it said to choose your own using a√x+b+c=d

OpenStudy (stefrheart):

Ah I understand hold for a minute please

OpenStudy (anonymous):

okay.

OpenStudy (stefrheart):

I dont think it has an extraneous solution because I got x=16

OpenStudy (anonymous):

well could you help me find one with one please? using the a√x+b+c=d formula thingy

OpenStudy (stefrheart):

Sure!

OpenStudy (stefrheart):

So you do understand what an extraneous solution is correct?

OpenStudy (anonymous):

i think it is how it solves the problem but if you plug it back in it doesnt work out

OpenStudy (anonymous):

like if you solve the radical but if you plug it in to check the answer will not work out

OpenStudy (stefrheart):

Yes that is correct! :)

OpenStudy (anonymous):

im not sure how to like create and solve a radical like that though... ;(

OpenStudy (anonymous):

Create two radical equations: one that has an extraneous solution, and one that does not have an extraneous solution. Use the equation below as a model. a√x+b+c=d Use a constant in place of each variable a, b, c, and d. You can use positive and negative constants in your equation.

OpenStudy (stefrheart):

is it \[a \sqrt{x+b+c}=d\] or \[a \sqrt{x} + b + c=d\]

OpenStudy (phi):

try something like \[ \sqrt{x+2} + 1 = -1 \]

OpenStudy (anonymous):

the first one @stefrheart

OpenStudy (anonymous):

@phi the only problem is is that i don't know how to solve that equation to check if it has an extraneous solution

OpenStudy (stefrheart):

To solve that all you need to do is subtract the one on both sides to create \[\sqrt{x+2}=-2\] and then square both sides \[(\sqrt{x+2})^{2}=(-2)^{}\] and then solve

OpenStudy (stefrheart):

just for phi's equation

OpenStudy (dinamix):

x=-8 @stefrheart

OpenStudy (dinamix):

@kenzi041

OpenStudy (anonymous):

i dont even know. i never learned radicals

OpenStudy (stefrheart):

oh you dont know what a radical is or how to work with it?

OpenStudy (anonymous):

we had a large learning gap because they changed curriculums

OpenStudy (anonymous):

not really.

OpenStudy (stefrheart):

Ah okay so you should learn what a radical is first then, before proceeding with the problem

OpenStudy (anonymous):

oh.

OpenStudy (stefrheart):

It would make alot more sense before solving this. Do you know what a \[x ^{2}\] is? It would be easier to teach you if you knew what it is

OpenStudy (anonymous):

x squared?

OpenStudy (stefrheart):

yes you know what to do with it if someone gave you the equation \[4^{2}\] ?

OpenStudy (anonymous):

16, i know like how to do that i just dont understand how to use like larger equations like a√x+b+c=d

OpenStudy (stefrheart):

Ah okay so you know how to do that. Well the larger equations are just the same ting as a regular square root.

OpenStudy (anonymous):

okay, so how do i find one with an extraneous solution

OpenStudy (anonymous):

because i have one for without but i just need one with one

OpenStudy (stefrheart):

I think I may have one hold on

OpenStudy (anonymous):

ok

OpenStudy (dinamix):

x=-8 u have check in the equation ans see it

OpenStudy (stefrheart):

Whats the extraneous solution then? @dinamix

OpenStudy (anonymous):

like how do you plug it in to check if it has one @dinamix

OpenStudy (anonymous):

i already tried but it was in a different format than what the question asked

OpenStudy (stefrheart):

Yeah I noticed ​​

OpenStudy (anonymous):

and i got more confused..

OpenStudy (stefrheart):

Hmmm im trying to figure out equations because in order for it to have both an actual and extraneous there needs to be 2 solutions which is a \[x ^{2}+bx+c\] unless we can do that you cant get both an actual and extraneous

OpenStudy (anonymous):

no it needs to just be a extraneous solution.. you write two problems and one should be regular and one should have an extraneous solution

OpenStudy (stefrheart):

Ohhh then I already have one then

OpenStudy (stefrheart):

\[-5\sqrt{-3x-5}=-25\]

OpenStudy (anonymous):

but doesnt it have to be x+b+c=d

OpenStudy (stefrheart):

if you wwant it you can just put -3x+-5+0 for under the radical

OpenStudy (anonymous):

okay, one sec im checking a problem i just did.

OpenStudy (stefrheart):

I will be back in a few minutes

OpenStudy (anonymous):

i think i found one, the problem i chose originally

OpenStudy (anonymous):

and okay

OpenStudy (stefrheart):

AH okay

OpenStudy (stefrheart):

Message me if need you still need help

OpenStudy (anonymous):

thanks for helping

OpenStudy (anonymous):

how do i give u a medal?

OpenStudy (stefrheart):

@kenzi041 click the best response

OpenStudy (anonymous):

@stefrheart i failed, but thanks anyway

OpenStudy (stefrheart):

Aw im sorry :( @kenzi041